A sample of dry air, with a volume of 125 ml and a temperature of 53°C, is heated at constant pressure to 95°C. What is the new volume in milliliters of the sample?

1
Expert's answer
2021-11-17T10:02:04-0500

General gas law:

(P1*V1) / T1 =(P2*V2)/ T2

Since p = const: V1/T1 = V2/T2

So, step 1. Convert the temperatures to Kelvin. {K=C+273}

T1 = 53 + 273 = 326 K

T2 = 95 + 273 = 368 K

(125 mL / 326 K) = (V2 mL / 368 K)

V2 = 141mL